ova Company’s total overhead cost at various levels of activity are presented below: Month Machine- Hours Total Overhead Cost April 49,000 $ 199,830 May 39,000 $ 173,130 June 59,000 $ 226,530 July 69,000 $ 253,230 Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 39,000 machine-hour level of activity is: Utilities (variable) $ 58,500 Supervisory salaries (fixed) 55,000 Maintenance (mixed) 59,630 Total overhead cost $ 173,130 Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements. Required: 1. Estimate how much of the $253,230 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $253,230 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) 2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX. 3. Express the company’s total overhead cost in the form Y = a + bX. 4. What total overhead cost would you expect to be incurred at an activity level of 44,000 machine-hours?
Answer #1
1) Utilities (58500/39000)*69000 103500
Supervisory Salaries 55,000
Maintenance (balancing figure) 94,730
Total overhead cost 253,230
Maintenance cost in july 94,730
2) Estimating formula for maintenance cot
MH's cost
high 69,000 94,730
low 39,000 59,630
Difference 30,000 35,100
Variable cost per unit = 35100/30000
1.17
Fixed cost = 947330 - 69000*1.17
14000
Cost formula
Y= 14,000 + 1.17 X
3) Companys total overhead cost
Variable Fixed
Utilities 1.5
Supervisory salaries 55,000
Maintenance cost 1.17 14,000
total 2.67 69000
cost formula
Y= 69,000 + 2.67 X
4) total overhead cost
Y= 69,000 + 2.67 X
so
fixed cost 69,000
Variable cost (44000*2.67) 117480
total overhead cost 186,480 answer